OpenStudy (anonymous):

Let F be a vector field F = (3x)i + (xz^3)j - (e^y)k. Let B be a solid ball x^2 + y^2 + z^2 <= 1. If n is the outward facing unit normal to S, use the divergence theorem to calculate the surface integral

OpenStudy (anonymous):

I know that you can convert it to the triple integral of del F dV.

OpenStudy (anonymous):

We have a solid \(B\) bounded by a surface \(S\).

OpenStudy (anonymous):

We will let \(\mathbf S\) parametrize \(S\).\[ \iint_S\mathbf F \times d\mathbf S =\iint_S\mathbf F\cdot \mathbf n~dS= \iiint_V\nabla \cdot \mathbf F~dV \]

OpenStudy (anonymous):

So we want to gets \(B\)'s limits of integration: \[\begin{array}{rcl} -\sqrt{1-(y^2+x^2)}<&z&<\sqrt{1-(y^2+x^2)}\\ -\sqrt{1-x^2}<&y&<\sqrt{1-x^2} \\ -1 < &x& < 1 \end{array} \]

OpenStudy (anonymous):

Then \(dV = dzdydx\).

OpenStudy (anonymous):

\[ \nabla \cdot F = 3 + 0 - 0 = 3 \]

OpenStudy (anonymous):

The only think we can do to make this any easier is to use spherical coordinates.

OpenStudy (anonymous):

Gotcha. I didn't quite understand the limits, but I am familiar with how you set it up.

OpenStudy (anonymous):

I would convert to spherical coords. I don't remember exactly certain details.

OpenStudy (anonymous):

I believe the Jacobian for it is \(dz~dy~dz = \rho \sin \phi ~d\rho ~d\phi ~d\theta \)

OpenStudy (anonymous):

rho^2 I believe

OpenStudy (anonymous):

Well, who cares anyway: \[ \iiint_V 3~dV = 3 \iiint dV = 3V = 3\frac 43\pi (1)^3 \]

OpenStudy (anonymous):

That was my other question. In all instances, if given that the object is of a known volume, you can substitute the triple integral of dV for the known formula, correct?

OpenStudy (anonymous):

And since we calculate out del F to be 3, its just substitute in for r and multiply by 3, as you did.

OpenStudy (anonymous):

No, you are assuming that \(\nabla \cdot \mathbf F\) is constant.

OpenStudy (anonymous):

Only if del F is a constant*

OpenStudy (anonymous):

Otherwise, you would integrate over the limits of \[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi}\int\limits_{0}^{1} delop F \rho^2 \sin \phi d \rho d \phi d \theta \]

OpenStudy (anonymous):

Given that this is a sphere.

OpenStudy (anonymous):

yeah, I guess. So \(0<\phi <\pi\) then? I forgot.

OpenStudy (anonymous):

For a sphere, yes.

OpenStudy (anonymous):

Sweet, thanks again

OpenStudy (anonymous):

If you happen to be bored, I'm asking one last question that concerns Stokes theorem. I've tried it a couple of times and I'm not sure where I'm going wrong exactly.

OpenStudy (anonymous):

ok